OpenStudy (anonymous):

Find the area of the triangle with M< A= 80°, b = 6m, c =4m. Round your answer to the nearest tenth if necessary. A. 12 m2 B. 23.6 m2 C. 11.8 m2 D. none of these

OpenStudy (raden):

area = 1/2 * b * c * sinA area = 1/2 * 6 * 4 * sin80 and the calculator

OpenStudy (anonymous):

So its C. thanks!!